# From First Patch to Long-Term Contributor: Evaluating Onboarding Recommendations for OSS Newcomers

### v1

@@ 1 Introduction
For an aspiring OSS candidate, having the first patch accepted is a significant milestone towards this goal. The sense of accomplishment and recognition of the first accepted patch motivates further contributions [7, 8, 9]. Several studies [10, 11, 12] have focused specifically on prospective OSS contributors who have not yet had changes accepted [5, 13]. This group is often called ‘newcomers’  [5, 13], and studies have proposed recommendations to increase newcomers’ chances of successful onboarding.
However, there are two uncertainties regarding the recommendations based on current research. First, most of those recommendations are based on a limited number of projects, which raises concerns about their generalizability across the diverse OSS spectrum. Let’s consider the scenario where a recommendation provides contradicting results in a different context. Following this recommendation in an unhelpful context may hinder a newcomer’s onboarding instead of helping. Second, it remains unclear whether the same recommendations also apply to non-newcomers. If some recommendations are specific to newcomers, contributors who continue to follow them after their first changes are accepted could harm their chances of becoming long-term contributors. Therefore, this study addresses these two research gaps by evaluating newcomers’ onboarding recommendations i) across a diverse group of OSS projects to assess whether they are context-dependent and ii) on non-newcomers to assess whether they are newcomer-specific.
@@
Motivation: Newcomer onboarding is essential for integrating new ideas and ensuring the sustainability of OSS projects. Due to its importance, prior studies have identified several recommendations to assist the onboarding of OSS newcomers [13, 14, 15, 16]. While our primary motivation behind this question is to empirically identify the recommendations and validate them in our RQ2, answering this question will also aggregate these recommendations into a common list that a newcomer can examine.
Method: Following the guidelines of Kitchenham and Carter [17], we conducted a Systematic Literature Review (SLR) that identified 14 primary studies.
Key results: We have identified 15 task-related recommendations for prospective newcomers to OSS projects, including nine recommendations on task selection, four on task preparation, and two regarding task completion.
@@
Key results: The results of our regression models suggest that four identified recommendations have significant positive associations with newcomers’ first patch acceptance in most contexts. Four recommendations are context-dependent, four indicate significant negative associations in most projects, and the remaining ones are either conclusive or could not be evaluated due to multicollinearity (defined in Section 3.2).
@@
Method: Using the same set of metrics and regression modeling strategy as RQ2, we evaluate compliance of those recommendations against the odds of a non-newcomers’ patch acceptance.
Key results: Our results found three newcomer-specific recommendations contradicting their associations between newcomers and non-newcomers. Therefore, once a contributor becomes a non-newcomer, they may change their task approaches to increase their odds of becoming long-term contributors.
@@
  • We have released our dataset and scripts publicly available on GitHub: https://github.com/WSU-SEAL/newcomer-onboarding
Organization: The remainder of the paper is organized as follows. Section 2 briefly overviews the related works. Section 3 presents the research methodology. Section 4 presents this study’s results. Section 5 discusses key implications based on our results. Section 6 addresses threats to validity for our findings. Finally, Section 7 concludes this paper.
@@ 2 Related Works
Challenges faced by newcomers: A newcomer’s journey to becoming a long-term contributor is full of challenges [5, 19, 20, 21, 22, 23, 24]. Steinmacher et al. identified 58 social barriers newcomers face for contributing their first change [5]. Besides social ones, newcomers also encounter technical challenges in contributing to large, matured OSS projects [20]. Prior studies also studied barriers faced by one-time contributors in OSS projects, who had exactly one patch accepted [25], and casual contributors who contribute occasionally [21]. These studies conclude that the primary challenges encountered by prospective OSS joiners include lack of support materials [25, 9, 6, 26], a steep learning curve [20, 19, 6], unfriendly project maintainers [6, 25, 27], and lack of timely feedback [28, 25, 6].
@@
Recommendations to assist newcomers’ onboarding: Several existing studies have identified recommendations to assist newcomers’ onboarding process. These studies have focused on various aspects that a newcomer can follow. Several studies have focused on Task selection, i.e., which task a newcomer should start with [13, 20, 29, 30, 31]. Some studies have also focused on Task preparation, i.e., how to prepare the changeset to facilitate easy reviewing from maintainers [32, 10]. Task completion-related suggestions focus on communicating and responding to reviewers’ suggestions after a change is submitted for reviews [27, 33, 20]. Socialization and communication-related guidelines focus on communication etiquette and strategies to connect with existing members  [12, 33, 26, 34]. Besides newcomers themselves, project maintainers also play important roles in fostering a supportive environment for newcomers [35]. Recommendations for maintainers include mentoring [14], creating a documentation outlining step-by-step onboarding process [36, 33, 37], ensuring that newcomers feel welcome [36], having and enforcing code of conduct [36, 38, 33], implementing automated task recommendation system [32], using specialized tags for tasks that newcomers can work on [39], using a community health dashboard to track newcomers’ progress [40], and creating a portal to help newcomers overcome common entry barriers [33].
@@
Automated tools for assisting newcomers Proper mentoring throughout the onboarding journey is essential to ensure newcomers’ future success [41, 42, 43]. To assist the newcomers in this journey, researchers have proposed an automated model to identify potential metors [44], conversational bots that can act as a mentor [45, 46] and answer newcomers’ questions [47]. Another direction focused on building models to automatically identify suitable tasks or bugs that newcomers can start with [48, 32, 49, 50, 51, 29, 39]. Since many OSS projects have steep learning curves [6], researchers have proposed tools to help newcomers gain knowledge of project goals and design by building project-specific knowledge-based project memory [52, 53], a protocol for exchanging information via integrated development environment [54], and reverse-engineered visualization tools [55].
@@
Novelty of this study Our study differs from prior works due to three novel analyses: First, we aim to provide a quantitative assessment of recommendations identified based on qualitative studies on a wide range of OSS projects. Second, we aim to identify if context dependence applies to any recommendations. Third, we aim to assess if those recommendations also apply to non-newcomers to assist the long-term success of OSS joiners.
@@
Tan et al.’s study to investigate the characteristics of Good First Issues (GFI) on GitHub is closest to this study. Their investigation focuses on the resolution process, contributors, textual characteristics, and labelers’ reputations, which are more likely to make a GFI effective (i.e., indeed solved by a newcomer). Our goal is vastly different as we focus on any change newcomers work on, including GFIs and non-GFIs, and how their task practices influence code acceptance. Due to minor conceptual similarity, five of their textual metrics also belong among the 15 selected in our study.
3 Research Method
@@
The following subsections detail our systematic literature review (SLR) methodology to identify a set of task-related onboarding recommendations and our empirical investigation to assess whether they are positively associated with newcomers’ and non-newcomers’ code acceptance.
@@ 3.1 SLR Methodology
We conducted an SLR of existing studies on OSS newcomers’ onboarding to identify task-related recommendations that a prospective newcomer can follow. We found two existing SLRs focusing on newcomers’ onboarding  [9, 6]. Since several recent studies have recommended new strategies to assist newcomers’ onboarding [32, 10, 13, 12], we decided to conduct a new SLR. Our SLR specifically focuses on finding recommendations regarding newcomers’ task selection and work practices, as those are crucial factors behind a newcomer’s successful onboarding [12]. We do not focus on recommendations that are targeted toward project maintainers. We designed our SLR based on the guidelines suggested by Kitchenham and Charters [17], and it has been the gold standard for conducting SLR in the SE domain. Table I provides an overview of the number of papers surviving during each of our SLR steps, as follows.
@@
Searching Google Scholar, IEEE, ACM, and Scopus databases and filtering based on titles 158
After removing duplicates 100
Study selection based on abstract 29
Study selection by a quick reading of the publications 18
Selected studies after Snowballing 26
Selected studies considering expert suggestions 31
Studies providing task-related recommendations 14
@@
Our primary goal of performing the SLR is to identify recommendations regarding newcomers’ task selection and work practices. For this goal, we defined four inclusion criteria (IC) and eight exclusion criteria (EC). Our inclusion criteria include:
@@
  • IC2: The publication provides recommendation(s) to newcomers of OSS projects regarding work practice such as task preparation, task presentation, and communication.
@@
We adopted a hybrid approach for performing SLR, which involves searching in Google Scholar, Scopus, IEEE, and ACM. Following the suggestions of prior studies [5, 9] and considering our final goal, we defined our search strings, using three categories of keywords, which include i) target audience, ii) goal, and iii) context. We have used the ‘advanced search’ option of the above-mentioned search engines. Table  II presents the three categories and lists the keywords belonging to each. While the search queries vary among the engines, the queries were formulated in such a way that each result must include at least one keyword from each of the three categories. A simplified search string looks like the following.
@@
The number of returned results for some search engines (e.g., Google Scholar) was large. As most of the results are irrelevant, it would have been extremely time-consuming to inspect all the results. Therefore, we went through the results pages containing ten entries individually. If the title of a paper seemed relevant, we stored the title and list of authors in a spreadsheet for further inspection. We stopped inspecting further results if five consecutive pages returned no relevant titles. We conducted database searches during the last week of August 2022 and identified 158 papers with potentially relevant titles. After removing the duplicate entries, we were left with 100 papers.
@@
In this step, we downloaded and read the abstracts for those remaining papers to exclude irrelevant ones. This step resulted in the exclusion of a total of 71 papers. In the next iteration, we skimmed through each of the remaining 29 papers to apply our first five exclusion criteria (i.e., EC1 -EC5). After discarding 11 papers based on EC1-EC5, we identified 18 studies at the end of this step.
@@
Backward Snowballing is the process of examining the references of a publication to select additional relevant studies. Using Backward Snowballing, we identified six additional studies. Forward Snowballing is identifying relevant publications that cited our selected preliminary studies. In this step, we found two additional relevant studies. At the end of this step, we found 26 studies.
@@
In this step, we decided to validate our list of 26 papers based on the opinions of expert researchers. We selected three expert researchers whose names appear in more than three studies. We sent an email to the three expert researchers mentioning the goal of this SLR and providing the list of 26 papers. We asked them whether they could suggest additional studies relevant to our goal. Two of the three expert researchers replied to our email, and we included five more studies in this step.
@@
In this step, we carefully reviewed the 31 publications and applied EC6, EC7, and EC8 (i.e., we discarded studies that did not provide recommendations regarding OSS newcomers’ task selection or work practices). A total of 14 studies survived this step. While adding five studies in Step 5 based on expert recommendation may question the adequacy of our search strings, we noticed that none of those five studies survived EC6, EC7, and EC8. Our investigation also revealed the lack of ‘goal’ related keywords as the primary reason behind their exclusion by our searches. Our replication package lists the discarded papers with deciding steps [56].
@@
In this step, two authors reviewed the 14 papers independently and highlighted sentences suggesting recommendations to newcomers about task selection or work practices. The disagreements between the two authors in identifying the recommendations were resolved in a discussion session. We compiled a list of the highlighted sentences in a separate document. We followed a card sorting technique suggested by Zimmermann [57] to identify the list of recommendations. We also mapped each of the recommendations to one or more primary studies.
3.2 Empirical Study Methodology
We followed a six-step method to answer RQ2 and RQ3. In the first step, we defined one quantitative metric to measure a code change’s compliance against each recommendation listed in Table VI. Second, we identified a set of projects to test each recommendation-based hypothesis (i.e., whether a recommendation has a positive or negative association with code acceptance). Third, we mine code review /pull request datasets for the selected projects. Fourth, we computed the attributes identified in the first step using our dataset. Fifth, we trained and evaluated logistic regression models for each project /project group following Harrell Jr.’s guidelines [18]. We trained two models for each dataset, one for newcomers (RQ2) and the other for non-newcomers (RQ3). We test each attribute’s significance using the Wald Chi-Square statistics (Wald $\chi^{2}$) test and direction of influence using the Odds Ratio (OR). Finally, we map the results of our regression models back to our recommendations to provide verdicts on whether our analyses support, contradict or are inconclusive in each recommendation. The following subsections detail these six steps.
@@ 3.2 Empirical Study Methodology
Table III: Patch characteristics-related recommendations obtained from Systematic Literature Review

### v2

@@ 1 Introduction
For an aspiring OSS candidate, having the first patch accepted is a significant milestone towards this goal. The sense of accomplishment and recognition of the first accepted patch motivates further contributions [7, 8, 9]. Several studies [10, 11, 12] have focused on prospective OSS contributors who have not yet had changes accepted  [5, 13]. According to Steinmacher et al., a person is considered a ‘newcomer’ to an OSS project until he/she gets his/her first code review/pull request accepted to that project [5]. Studies have proposed recommendations to increase newcomers’ chances of successful onboarding.
However, there are two uncertainties regarding the recommendations based on current research. First, due to the diverse nature of OSS, it’s uncertain how well these recommendations will apply to other OSS projects. Let’s consider the scenario where a recommendation provides contradicting results in a different context. Following this recommendation in an unhelpful context may hinder a newcomer’s onboarding instead of helping. Second, it remains unclear whether the same recommendations also apply to non-newcomers. If some recommendations are specific to newcomers, contributors who continue to follow them after their first changes are accepted could harm their chances of becoming long-term contributors. Therefore, this study addresses these two research gaps by evaluating newcomers’ onboarding recommendations i) across a diverse group of OSS projects to assess whether they are context-dependent and ii) on non-newcomers to assess whether they are newcomer-specific.
@@
Motivation: Newcomer onboarding is essential for integrating new ideas and ensuring the sustainability of OSS projects. Due to its importance, prior studies have identified several recommendations to assist the onboarding of OSS newcomers [13, 14, 15, 16]. While our primary motivation behind this question is to identify the recommendations and validate them in our RQ2 empirically, answering this question will also aggregate these recommendations into a common list that a newcomer can examine.
Method: Following the guidelines of Kitchenham et al. [17], we conducted a Systematic Literature Review (SLR) that identified 16 primary studies.
Key results: We have identified 15 task-related recommendations for prospective newcomers to OSS projects, which include nine recommendations on task selection, four on task preparation, and two regarding task completion.
@@
Key results: The results of our regression models suggest that four identified recommendations have significant positive associations with newcomers’ first patch acceptance in most contexts. Four recommendations are context-dependent, four indicate significant negative associations in most projects, and the remaining ones are either inconclusive or could not be evaluated due to multicollinearity (defined in Section 4.1.5).
@@
Method: Using the same set of metrics and regression modeling strategy as RQ2, we evaluate compliance of those recommendations against the odds of a non-newcomer’s patch acceptance.
Key results: Our results found three newcomer-specific recommendations contradicting their associations between newcomers and non-newcomers. Therefore, OSS candidates should avoid using these three strategies once their first patches are accepted to enhance their chances of becoming long-term contributors.
@@
  • We have released our dataset and scripts on GitHub [^1] https://github.com/WSU-SEAL/newcomer-onboarding.
Organization: The remainder of the paper is organized as follows. Section 2 briefly overviews the related works. Section 3 details the methodology and results of our first stage (i.e., SLR). Section 4 details the methodology and results of empirical evaluation to investigate RQ2 and RQ3. Section 5 discusses key implications based on our results. Section 6 addresses threats to validity for our findings. Finally, Section 7 concludes this paper.
@@ 2 Related Works
Challenges faced by newcomers: A newcomer’s journey to becoming a long-term contributor is full of challenges [5, 19, 20, 21, 22, 23, 24]. Steinmacher et al. identified 58 social barriers newcomers face for contributing their first change [5]. Besides social ones, newcomers also encounter technical challenges in contributing to large, matured OSS projects [20]. Prior studies also studied barriers faced by one-time contributors in OSS projects, who had exactly one patch accepted [25],[26], quasi-contributors (newcomers without any accepted CR/PR)[27], and casual contributors who contribute occasionally [21]. Another couple of studies focused on drive-by commits [28] and episodic volunteering[29]. These studies conclude that the primary challenges encountered by prospective OSS joiners include lack of support materials [25, 9, 6, 30], a steep learning curve [20, 19, 6], unfriendly project maintainers [6, 25, 31], and lack of timely feedback [32, 25, 6].
@@
Recommendations to assist newcomers’ onboarding: Several existing studies have identified recommendations to assist newcomers’ onboarding process. These studies have focused on various aspects that a newcomer can follow. Studies have focused on Task selection, i.e., which task a newcomer should start with [13, 20, 33, 34, 35]. Some studies have also focused on Task preparation, i.e., how to prepare the changeset to facilitate easy reviewing from maintainers [36, 10]. Task completion-related suggestions focus on communicating and responding to reviewers’ suggestions after a change is submitted for reviews [31, 37, 20]. Socialization and communication-related guidelines focus on etiquette and strategies to connect with existing members  [12, 37, 30, 38]. Besides newcomers themselves, project maintainers also play important roles in fostering a supportive environment for newcomers [39]. Recommendations for maintainers include mentoring [14], creating a documentation outlining step-by-step onboarding process [40, 37, 41], ensuring that newcomers feel welcome [40], having and enforcing code of conduct [40, 42, 37], implementing automated task recommendation system [36], using specialized tags for tasks that newcomers can work on [43], using a community health dashboard to track newcomers’ progress [44], and creating a portal to help newcomers overcome common entry barriers [37].
@@
Automated tools for assisting newcomers: Proper mentoring throughout the onboarding journey is essential to ensure newcomers’ future success [45, 46, 47]. To assist the newcomers in this journey, researchers have proposed an automated model to identify potential mentors [48], conversational bots that can act as a mentor [49, 50] and answer newcomers’ questions [51]. Another direction focused on building models to automatically identify suitable tasks or bugs that newcomers can start with [52, 36, 53, 54, 55, 33, 43]. Since many OSS projects have steep learning curves [6], researchers have proposed tools to help newcomers gain knowledge of project goals and design by building project-specific knowledge-based project memory [56, 57], a protocol for exchanging information via integrated development environment [58], and reverse-engineered visualization tools [59].
@@
Novelty of this study: Our study differs from prior works based on three novelties. First, we quantitatively assess all recommendations on a wide range of OSS projects. Such assessments are crucial since many recommendations are based on qualitative investigations. Second, we investigate if context dependence applies to any of these recommendations. Third, we aim to assess if those recommendations also apply to non-newcomers to assist the long-term task strategies of OSS candidates.
@@
Closely related work: Tan et al.’s study to investigate the characteristics of Good First Issues (GFI) on GitHub [10] is the closest one to ours. They examined correlations between a GFI effectiveness (i.e., indeed solved by a newcomer) and its resolution process, contributors’ characteristics, textual characteristics, and labelers’ reputations. As Tan et al.’s characterization of effective GFIs overlaps with this study’s Task Selection aspect, five of their textual metrics also belong among the 15 selected in our study. However, this study differs from Tan et al. [10] in the following three key areas. First, this study aims to help newcomers select, prepare, and complete tasks. On the other hand, Tan et al. [10] aims to help maintainers label more effective GFIs. Second, we examine any change newcomers worked on, including GFIs and non-GFIs. Finally, we also empirically evaluate how newcomers’ and non-newcomers’ task preparation and completion practices influence code acceptance.
3 Stage 1: Identification of task-related recommendations for newcomers
@@
The following subsections detail our systematic literature review (SLR) methodology to identify a set of task-related onboarding recommendations and SLR results.
@@ 3.1 SLR Methodology
We conducted an SLR of existing studies on OSS newcomers’ onboarding to identify task-related recommendations that a prospective newcomer can follow. We found two existing SLRs focusing on newcomers’ barriers [9, 6]. Since the previous SLRs were not conducted in the context of newcomers’ task-related recommendations and several recent studies have recommended new strategies to assist newcomers’ onboarding [36, 10, 13, 12], we decided to conduct a new SLR. Our SLR specifically focuses on finding recommendations regarding newcomers’ task selection and work practices, as those are crucial factors behind a newcomer’s successful onboarding [12]. We do not focus on recommendations that are targeted toward project maintainers. We designed our SLR based on the guidelines suggested by Kitchenham et al. [17], and it has been the gold standard for conducting SLR in the SE domain. Table I provides the number of papers surviving each SLR step. The steps are as follows.
@@
Searching Google Scholar, IEEE, ACM, and Scopus databases 605
After removing duplicates 498
Study selection based on title and abstract 68
Study selection by a quick reading of the publications 32
Inclusion of additional studies based on Snowballing 40
After adding studies based on expert suggestions 45
Studies surviving all eight exclusion criteria 16
@@
Our primary goal of performing the SLR is to identify recommendations regarding newcomers’ task selection and work practices. For this goal, we defined four inclusion criteria (IC) and eight exclusion criteria (EC). Our inclusion criteria are:
@@
  • IC2: The publication provides recommendation(s) to newcomers to OSS projects regarding work practice such as task preparation, task presentation, and communication.
@@
We adopted a hybrid approach for performing SLR, which involves searching in Google Scholar, Scopus, IEEE, and ACM. Following the suggestions of prior studies [5, 9] and considering our final goal, we defined our search strings using three categories of keywords, which include i) target audience, ii) goal, and iii) context. We have used the ‘advanced search’ option of the above-mentioned search engines. Table  II presents the three categories and lists the keywords belonging to each. While the search queries vary among the engines, the queries were formulated in such a way that each result must include at least one keyword from each of the three categories. A simplified search string looks like the following.
@@
We noticed that the number of entries returned by Google Scholar was large. As most of the results are irrelevant, inspecting all entries would have been extremely time-consuming. Therefore, we went through the results pages containing ten entries individually. If the title of a paper seemed relevant, we stored the title and list of authors in a spreadsheet for further inspection. We stopped inspecting further results if five consecutive pages returned no relevant titles, as Google Scholar ranks search results by the relevance of search query [60]. We conducted two database searches. Our first search was conducted during the last week of August 2022. In the first search, we selected papers based on study titles from the results. However, this protocol is not fully compliant with Kitchenham et al.’s guidelines [17], which recommend inspecting both titles and abstracts before excluding any search result. Therefore, we conducted a fully compliant second search during the second week of October 2024. Moreover, this second search also helped us include papers published since our first search. Our database search found 605 papers, of which 349 were found in the ACM database, 49 were from the Google Scholar database, 28 were from the IEEE database, and 179 were from the Scopus. After duplicate removal, we were left with 498 studies.
@@
In this step, we examined the titles and abstracts of the 498 identified studies and excluded 430 that we deemed irrelevant to our SLR goals. We then applied our five exclusion criteria (i.e., EC1-EC5) to the remaining 68 studies to exclude 36. The number of studies filtered based on EC1-EC5 is 8, 6, 3, 12, and 7, respectively.
@@
Backward Snowballing is the process of examining the references of a publication to select additional relevant studies. Using Backward Snowballing, we identified six additional studies. Forward Snowballing is identifying relevant publications that cited our selected preliminary studies. In this step, we found two additional relevant studies. At the end of this step, we had 40 studies, among which eight came from the snowballing steps.
@@
In this step, we decided to validate our list of 40 studies based on the opinions of expert researchers. We selected three expert researchers whose names appear in more than three studies. We sent them emails mentioning this SLR’s goal and providing the list of 40 papers. We explicitly solicited their suggestions regarding studies missing from our list. Two of the three expert researchers replied to our email, and we included five more studies based on their suggestions.
@@
In this step, we carefully reviewed the 45 publications and applied EC6, EC7, and EC8 (i.e., we discarded studies that did not provide recommendations regarding OSS newcomers’ task selection or work practices). A total of 16 studies survived this step. While adding five studies in Step 5 based on expert recommendation may question the adequacy of our search strings, we noticed that none of those five studies survived EC6, EC7, and EC8. Our investigation also revealed the lack of ‘goal’ related keywords as the primary reason behind their exclusion by our searches. Our replication package lists the discarded papers with deciding steps [61]. Among the excluded 29 studies, 2 were due to EC7 and the remaining 27 were due to EC8.
@@
In this step, two authors reviewed the 16 papers independently and highlighted sentences suggesting recommendations to newcomers about task selection or work practices. The disagreements between the two authors in identifying the recommendations were resolved in a discussion session. We compiled a list of the highlighted sentences in a separate document. We followed a card sorting technique suggested by Zimmermann [62] to identify the list of recommendations. We also mapped each of the recommendations to one or more primary studies.
Table III: Research methodology for studies on newcomers’ recommendations
Research method # Number of studies
Interview / survey 10 SLR1, SLR2, SLR4, SLR6, SLR7, SLR8, SLR9, SLR10, SLR11, SLR16
Mining software repositories (MSR) 5 SLR3, SLR12, SLR13, SLR14, SLR15
Mixed: Survey and MSR 1 SLR5
@@ 3.2 RQ1: What are the task-related recommendations for prospective newcomers of OSS projects?
Table IV: Task-related recommendations for prospective newcomers identified based on our Systematic Literature Review
#ID Recommendation Rationale Studies
Recommendation group: Task selection
R1 Work on smaller tasks first and later choose larger tasks Working on smaller tasks might increase the likelihood of newcomers’ first patch getting accepted SLR1[31], SLR2[14], SLR3[43], SLR5[15], SLR6[30], SLR8[37], SLR11[12], SLR12[13], SLR15[63]
R4 Start your contribution by working on minor bug fixes Minor bug fixes have lower knowledge requirements and also help improve the quality of a project SLR3[43], SLR5[15], SLR12[13], SLR14[36], SLR15[63]
R6 Start your contribution by performing less complex tasks Complex tasks are difficult for newcomers to solve and can possibly demotivate newcomers due to rejections SLR1[31], SLR3[43], SLR4[20], SLR7[34], SLR8[37], SLR9[64], SLR10[5], SLR13[10], SLR14[36], SLR16[65]
R7 Start your contribution by working on documentation changes Documentation changes are less complex, have fewer contribution barriers such as setting up environment or testing, and therefore have a higher odds of getting accepted SLR3[43], SLR5[15], SLR12[13], SLR14[36], SLR15[63]
Recommendation group: Task preparation
Recommendation group: Task completion
Table III shows an overview of the research method from the 16 selected primary studies. Ten out of the 16 studies are qualitative, and most of those recommendations did not include any quantitative evaluation of those recommendations. The five quantitative studies are recent works on characterizing GFIs on GitHub but lack analysis of non-GFIs. This result further shows the need for a large-scale quantitative analysis. Table IV shows our list of 15 recommendations identified based on our SLR. We divided the recommendations into three groups: Task selection, Task preparation, and Task completion. Under the task selection group, we got nine recommendations. The two most popular recommendations within this group for newcomers are to work on less complex tasks (R6), found in 9 primary studies, and to work on smaller tasks (R1), found in 8 studies. Within the Task preparation group, we obtained four recommendations, where making the title and description comprehensible ranks first with three studies. We found two Task completion recommendations (i.e., what to do after patch submission). While most recommendations are congruent, we found a conflict between R3 and R4, i.e., starting with new features vs. bug fixes.
4 Stage 2: Empirical evaluation of newcomers’ onboarding recommendations
The following subsections detail our research methodology to empirically evaluate the task-related recommendations identified based on our SLR, as shown in Table IV, and investigation results.
4.1 Research Method
We followed a six-step method to answer RQ2 and RQ3. In the first step, we defined one quantitative metric to measure a code change’s compliance against each recommendation listed in Table IV. Second, we identified a set of projects to test each recommendation-based hypothesis (i.e., whether a recommendation has a positive or negative association with code acceptance). Third, we mine code review /pull request datasets for the selected projects. Fourth, we computed the attributes identified in the first step using our dataset. Fifth, we trained and evaluated logistic regression models for each project /project group following Harrell Jr.’s guidelines [18]. We trained two models for each dataset, one for newcomers (RQ2) and the other for non-newcomers (RQ3). We test each attribute’s significance using the Wald Chi-Square statistics (Wald $\chi^{2}$) test and direction of influence using the Odds Ratio (OR). Finally, we map the results of our regression models back to our recommendations to provide verdicts on whether our analyses support, contradict, or are inconclusive in each recommendation. The following subsections detail these six steps.
Table V: Quantitative metrics selected for recommendations obtained based on SLR